Can someone verify the programming below using visual studio code python

  1. Stores 2 decimal numbers as variables.
  2. Stores one addition, one subtraction, one multiplication, and one division operation of these variables as variables.
  3. Prints out each of the four results as:

    <numeric value of variable 1> plus <numeric value of variable 2> equals <value of variable that stored the result of addition>

    <numeric value of variable 1> minus <numeric value of variable 2> equals <value of variable that stored the result of subtraction>

    ...etc. Each output should be on its own line.
  4. Stores the current month as a string variable (e.g. March, June, etc.) and day of the month as a numeric variable.
  5. Outputs "Today is day <day of month> of the month of <month variable>. This should be on a new line and tabbed over two times.
Answer #1


Given below is the code for the question. PLEASE MAKE SURE INDENTATION IS EXACTLY AS SHOWN IN IMAGE.
Please do rate the answer if it helped. Thank you.

num1 = 3.8
num2 = 1.5
sum = num1 + num2
diff = num1 - num2
prod = num1 * num2
quo = num1 / num2
print('{} plus {} equals {}'.format(num1, num2, sum))
print('{} minus {} equals {}'. format(num1, num2, diff))
print('{} multiplied by {} equals {}'.format(num1, num2, prod))
print('{} divided by {} equals {}'.format(num1, num2, quo))

month = 'July'
day = 30
print('\t\tToday is day {} of the month of {}'.format(day, month))

1 2 3 4 5 6 7 8 9 10 11 12 13 14 numl. = 3.8 num2 = 1.5 sum = numl. +. num2 diff = numl. -. num2 prod = numl. *. num2 quo = n

3.8 plus 1.5 equals 5.3 13.8 minus 1.5 equals 2.3 3.8 multiplied by 1.5 equals 5.699999999999999 13.8 divided by 1.5 equals 2